OpenStudy (natsent06):

PLEASE ANSWER FAST!! 21 is 35% of what number?

OpenStudy (acxbox22):

\[\frac{ 21 }{ x}=\frac{ 35 }{ 100}\] cross multiply 35x=2100 x=60

OpenStudy (natsent06):

Thank you!

OpenStudy (anonymous):

This question needs to be closed.

OpenStudy (natsent06):

ok 

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

what?

Elsa213 (elsa213):

Let the number be x Then 35% of x = 21 => 35/100 * x = 21 => x = 21*100/35 => x = 2100/35 = 60 Therefore, 35% of 60 = 21 Ans